# Integrable sub-Riemannian geodesic flows on the special orthogonal group

### v2

@@ 3.2. A Lax pair and bi-Hamiltonian structure
The bi-Hamiltonian structure we have constructed is a variation of that for the rigid body given in [27] (see also the construction in [5]) who are careful to note that if any of the $J_{i}$ are zero then this “…can also be employed for flat bodies, but the construction of the Hamiltonian functions is more complicated”. Indeed, we observe that the altered Hamiltonian $F$ in (17) differs from the form given in [27, Eq. 3.8]. Nevertheless, as confirmed in their concluding remarks, the proof that the integrals are in involution using a recursion relation may still be applied. Indeed, one can reproduce their arguments exactly to show that the sub-Riemannian Manakov integrals commute.
@@ 3.3. Proof of Theorem 1
The integrability of the sub-Riemannian geodesic system follows directly from an application of Mishchenko and Fomenko’s results which generalise the Euler equations of the rigid body to arbitrary semisimple Lie algebras.
Let $\mathfrak{g}$ be the complex semisimple Lie algebra $\mathfrak{gl}_{n}\mathbf{C}$ with ‘compact form’ $\mathfrak{g}_{u}=\mathfrak{u}(n)$ and the ‘normal form’ $\mathfrak{g}_{n}=\mathfrak{so}(n)$. One may verify the identity
(21) $[dH_{\text{sR}},A]=[M,B],$
from which the integrability of the equations of motion (13) follow directly from [24, Theorem 5.2]. More specifically, the integrals may be chosen from the collection of sub-Riemannian Manakov integrals, which are all of the form
$f(M+\sigma A),$
for $f$ a Casimir. In order to apply this theorem we note that although $A$ has a single zero-eigenvalue, it is still a regular element in $\mathfrak{gl}_{n}\mathbf{C}$ since all of its eigenvalues are distinct. The Lax pair formulation (15) also follows (see for instance the statement in [9, Theorem 1]).
@@
for $\hat{A}=\text{diag}(A_{n-k},\dots,A_{n})$. When interpreted as a rolling distribution this corresponds to the problem of rolling the (oriented) Grassmannian $\mathbb{Gr}_{k,n}^{+}$ along its affine tangent space. This problem has been considered extensively before in [17, 19] (see also the references therein) for the particular case of the standard Euclidean metric where $\hat{A}$ may be taken to be the idenity matrix. We remark however, that the result of Mishchenko and Fomenko used above does not work for $k>1$. The reason for this is that, although Eq. (21) holds for $A=\textrm{diag}(0,\dots,0,A_{n-k},\dots,A_{n})$ and $B=\textrm{diag}(0,\dots,0,1,\dots,1)$, the matrix $A$ is not regular, and so the proof in [24] no longer applies.
4. Further Work
@@
It is somewhat frustrating that we must assume the $A_{2},\dots,A_{n}$ to be distinct. This frustration is shared with the rigid body, where the integrability results of Manakov and Mishchenko & Fomenko both require the $J_{i}$ to be distinct. In the event that these become equal we obtain the so-called symmetric body. The proof of integrability breaks down as dependencies form between the Manakov integrals which were previously independent. On the other hand, the added symmetry of the body gives rise to additional conserved momenta. These, however, do not necessarily commute with one another, so one must consider the related notion of non-commutative integrability, as performed in [14, 15]. One could adapt their work to the sub-Riemannian case where some of the $A_{i}$s become equal.
Finally, we conclude with a conjecture. The famous Mishchenko-Fomenko conjecture (now a theorem [32], see also [8]) asked whether every Lie group admits an integrable Riemannian metric. This was intially proven by Mishchenko & Fomenko for semisimple Lie groups in [24]. Our work motivates a related sub-Riemannian extension of this question: for any (semisimple) Lie group $G$, does their exist an integrable sub-Riemannian metric of any given rank $k$? The known Riemannian case corresponds to $k=\dim G$. In this paper we have showed that for $G=\mathbf{SO}(n)$, the result is true for the distribution $\mathcal{D}$ with rank $k=n-1$. The general case would shed light on the class of systems discussed in [7] and [21].
@@ Acknowledgements
We wish to express our gratitude to the mathematics department at the University of Michigan for enabling this collaboration. We also thank Richard Montgomery for his helpful and constructive correspondence throughout the preparation of this paper. This research was supported in part by NSF grant 2103026 and AFOSR grants FA9550-23-1-0215 and FA9550-23-1-0400

### v3

@@ 3.2. A Lax pair and bi-Hamiltonian structure
The bi-Hamiltonian structure we have constructed is a variation of that for the rigid body given in [33] (see also the construction in [7]) who are careful to note that if any of the $J_{i}$ are zero then this “…can also be employed for flat bodies, but the construction of the Hamiltonian functions is more complicated”. Indeed, we observe that the altered Hamiltonian $F$ in (21) differs from the form given in [33, Eq. 3.8]. Nevertheless, as confirmed in their concluding remarks, the proof that the integrals are in involution using a recursion relation may still be applied. Indeed, one can reproduce their arguments exactly to show that the sub-Riemannian Manakov integrals commute.
3.3. The four-dimensional case
The case for $n=4$ is the smallest case to feature a non-trivial sub-Riemannian Manakov integral. The Hamiltonian from Eq (7) is
$H_{\textrm{sR}}(M)=\frac{1}{2}\left(\frac{M_{12}^{2}}{A_{2}}+\frac{M_{13}^{2}}{A_{3}}+\frac{M_{14}^{2}}{A_{4}}\right),$
for $M$ in $\mathfrak{so}(4)$. The 6-dimensional Lie algebra $\mathfrak{so}(4)$ possesses two Casimirs,
$C_{1}$ $=M_{12}^{2}+M_{13}^{2}+M_{14}^{2}+M_{34}^{2}+M_{24}^{2}+M_{23}^{2},$
$C_{2}$ $=M_{12}M_{34}-M_{13}M_{24}+M_{14}M_{23}.$
It is convenient to use the decomposition given in Eq. (15) together with the hat-map to identity $\widehat{m}$ in $\mathfrak{so}(3)$ with a vector $m$ in $\mathbf{R}^{3}$. In this way
$\mu=\begin{pmatrix}-M_{12}\\ -M_{13}\\ -M_{14}\end{pmatrix},\leavevmode\nobreak\ m=\begin{pmatrix}-M_{34}\\ M_{24}\\ -M_{23}\end{pmatrix}$,
and the Casimirs are compactly written
$C_{1}$ $=|\mu|^{2}+|m|^{2},$
$C_{2}$ $=\mu^{T}m.$
By considering the map $(\mu,m)\mapsto(\mu+m,\mu-m)$, we see that the set of constant $C_{1}$ and $C_{2}$ is diffeomorphic to the product of two spheres with respective radii $C_{1}+2C_{2}$ and $C_{1}-2C_{2}$. Incidentally, this establishes that the Casimirs satisfy
$C_{1}\geq 0,\text{ and }C_{1}^{2}\geq 4C_{2}^{2}.$
The regular coadjoint orbits are thus 4-dimensional. Integrability, therefore, requires one further integral in addition to the Hamiltonian. This is given by the sub-Riemannian integral $g_{3,2}$, which for convenience we scale by $\frac{1}{3}$ and denote by
$F(M)=A_{2}M_{12}^{2}+A_{3}M_{13}^{2}+A_{4}M_{14}^{2}+(A_{3}+A_{4})M_{34}^{2}+(A_{2}+A_{4})M_{24}^{2}+(M_{2}+M_{3})M_{23}^{2}.$
Hence, by Liouville’s theorem, $\mathfrak{so}(4)$ is foliated almost everywhere into tori equal to the connected components of the regular fibres of $(H_{\textrm{sR}},F,C_{1},C_{2})$.
Finally, we consider the analogous ‘rolling motion’ as in Section 2.4 for the case of a 4-dimensional ball rolling without spinning along a hyperplane. We proceed in greater generality, considering a unit ball in $\mathbf{R}^{n}$ rolling without slipping along the hyperplane $y_{1}=-1$.
Fix an orthonormal frame inside the centre of the ball which is coincident with the standard frame in space at time $t=0$. Let $c(t)$ be the vector for the point of contact between the ball and the hyperplane viewed from within the ball’s frame. For $t=0$ this is $c_{0}=(-1,0,\dots,0)^{T}$, and for later $t$ is given by $c(t)=g(t)^{-1}c_{0}$, where $g(t)$ is the matrix in $\mathbf{SO}(n)$ whose columns form the ball’s frame in space at time $t$. The velocity of the point of contact viewed from within the ball’s frame is
$\dot{c}=-g^{-1}\dot{g}g^{-1}c_{0}.$
To convert vectors from the ball’s frame to the space frame, we need only multiply on the left by $g$. Therefore, the velocity of the contact point in space is $\dot{g}g^{-1}c_{0}$, and thus, the no-slip condition implies that the point of contact satisfies
(25) $\frac{d}{dt}y_{j}=\Omega_{1j},\quad\text{for }2\leq j\leq n.$
Here we have introduced the right-invariant velocity $\Omega=\dot{g}g^{-1}$ in $\mathfrak{so}(n)$. The no-spin condition implies that $\Omega$ must belong to the rubber-rolling distribution given in Eq. (1), and the metric given in Eq. 6 defines the sub-Riemannian metric on $\mathcal{D}$ as in Eq. (2).
The equations of motion found in Eq. (17) define the following system in $\mathfrak{so}(4)$
$\begin{array}[]{ll}\displaystyle\frac{dM_{12}}{dt}=-\left(\frac{M_{13}M_{23}}{A_{3}}+\frac{M_{14}M_{24}}{A_{4}}\right),&\displaystyle\frac{dM_{23}}{dt}=\left(\frac{A_{2}-A_{3}}{A_{2}A_{3}}\right)M_{12}M_{13},\\ \displaystyle\frac{dM_{13}}{dt}=\left(\frac{M_{12}M_{23}}{A_{2}}-\frac{M_{14}M_{34}}{A_{4}}\right),&\displaystyle\frac{dM_{24}}{dt}=\left(\frac{A_{2}-A_{4}}{A_{2}A_{4}}\right)M_{12}M_{14},\\ \displaystyle\frac{dM_{14}}{dt}=\left(\frac{M_{12}M_{24}}{A_{2}}+\frac{M_{13}M_{34}}{A_{3}}\right),&\displaystyle\frac{dM_{34}}{dt}=\left(\frac{A_{3}-A_{4}}{A_{3}A_{4}}\right)M_{13}M_{14}.\end{array}$
The cometric relates the angular velocity with the angular momentum via $M_{1j}=A_{j}\Omega_{1j}$ for $2\leq j\leq n$. Therefore, we may find the motion for the point of contact of the ball by first integrating the system above, and then integrating Eq. (25). Examples of these curves for $n=4$ and for $(A_{2},A_{3},A_{4})=(2,3,4)$ are shown in Figures 2 and 3.
[…]
Figure 2. A bounded rolling curve.
[…]
Figure 3. An unbounded rolling curve.
@@ 3.4. Proof of Theorem 1
The integrability of the sub-Riemannian geodesic system follows directly from an application of Mishchenko and Fomenko’s generalisation of the Euler equations for the rigid body to arbitrary semisimple Lie algebras. For convenience we reproduce their result below [30, Theorem 5.2] (see also the statement in [11, Theorem 1]).
Theorem 2.
Let $\mathfrak{g}_{n}\subseteq\mathfrak{g}_{u}$ be the normal subalgebra of the semisimple complex algebra $\mathfrak{g}$ with compact real form $\mathfrak{g}_{u}$ and consider the Euler equations
$\frac{d}{dt}M=[M,\varphi(M)]$
defined on the subalgebra $\mathfrak{g}_{n}$ where $\varphi\colon\mathfrak{g}_{n}\rightarrow\mathfrak{g}_{n}$ is a self-adjoint linear map satisfying
(26) $\varphi(M),A]=[M,B]$
for some commuting $A,B\in\mathfrak{g}$ with $A$ regular. Then for any Casimir $f$ of $\mathfrak{g}_{n}$ the functions of the form $f(M+\sigma A)$ are constants of motion for the Euler equations and are in pairwise involution. Furthermore, among this set of functions, one can choose functionally independent integrals equal in number to half the dimension of a regular adjoint orbit of $\mathfrak{g}_{n}$.
To apply this theorem we set $\mathfrak{g}=\mathfrak{gl}_{n}\mathbf{C}$ with compact form $\mathfrak{g}_{u}=\mathfrak{u}(n)$ and normal form $\mathfrak{g}_{n}=\mathfrak{so}(n)$. The sectional operator $\varphi$ is identified with the map $M\mapsto dH_{\text{sR}}|_{M}$ and the identity (26) is satisfied as in (19) for $A=\text{diag}(0,A_{2},\dots,A_{n})$ and $B=\text{diag}(0,1,\dots,1)$. Casimirs are invariants of the adjoint representation, which for $\mathfrak{so}(n)$ are the trace-powers $f(M)=\operatorname{Tr}M^{k}$. Therefore, it follows from the theorem that among the sub-Riemannian Manakov integrals $g_{k,l}$ introduced in (20), there exist
$\frac{1}{2}\left(\dim\mathfrak{so}(n)-\operatorname{\textrm{rank}}\mathfrak{so}(n)\right)=\frac{1}{2}\left(\frac{n}{2}(n-1)-\left\lfloor\frac{n}{2}\right\rfloor\right)$
independent first integrals in involution, which defines an integrable system on the regular adjoint orbits of $\mathfrak{so}(n)$. This completes the proof of Theorem 1.∎
@@
for $\hat{A}=\text{diag}(A_{n-k},\dots,A_{n})$. When interpreted as a rolling distribution, this corresponds to the problem of rolling the (oriented) Grassmannian $\widetilde{\mathbf{Gr}}(k,n)$ along its affine tangent space. This problem has been considered extensively before in [20, 24] (see also the references therein) for the particular case of the standard Euclidean metric where $\hat{A}$ may be taken to be the identity matrix. We remark however, that the result of Mishchenko and Fomenko used above does not work for $k>1$. The reason for this is that, although Eq. (19) holds for $A=\textrm{diag}(0,\dots,0,A_{n-k},\dots,A_{n})$ and $B=\textrm{diag}(0,\dots,0,1,\dots,1)$, the matrix $A$ is not regular, and so the proof in [30] no longer applies.
Remark 3.4.
It is somewhat frustrating that we must assume the $A_{2},\dots,A_{n}$ to be distinct in order to ensure that $A$ remains regular. This frustration is shared with the rigid body, where the integrability results of Manakov, Mishchenko, and Fomenko both require the $J_{i}$ to be distinct. In the event that these become equal, we obtain the so-called symmetric body. The proof of integrability breaks down as dependencies form between the Manakov integrals which were previously independent. On the other hand, the added symmetry of the body gives rise to additional conserved momenta. These, however, do not necessarily commute with one another, so one must consider the related notion of non-commutative integrability, as performed in [16, 18]. One can adapt their work to the sub-Riemannian case where some of the $A_{i}$s become equal. A related work is [21] in which it is shown that specific sub-Riemannian metrics, which include our example for $A_{2}=\dots=A_{n}$, are integrable in the non-commutative sense.
Alternatively, there is a result due to Vinberg [41] which states that the limit of the algebra of integrals defined in Theorem 2 remains a commutative algebra with the same transcendence degree, even if $A$ tends to a non-regular element. This can be used to establish complete integrability of the system as various $A_{i}$s converge to each other. We remark that the resulting integrable system is larger than the algebra of integrals defined in Theorem 2 and depends on the way in which the regular $A$ converges to a non-regular element. These integrals and their dependence on the limit can be found explicitly by applying the work of Shuvalov [39].
4. Concluding Remarks
@@
Finally, we conclude with a conjecture. The famous Mishchenko-Fomenko conjecture (now a theorem [38], see also [10]) asked whether every Lie group admits an integrable Riemannian metric. This was initially proven by Mishchenko & Fomenko for semisimple Lie groups in [30]. Our work motivates a related sub-Riemannian extension of this question: for any (semisimple) Lie group $G$, does their exist an integrable sub-Riemannian metric of any given rank $k$? The known Riemannian case corresponds to $k=\dim G$. In this paper we have showed that for $G=\mathbf{SO}(n)$, the result is true for the distribution $\mathcal{D}$ with rank k=n1k=n-1. The general case would shed light on the class of systems discussed in [9] and [27].
@@ Acknowledgements
We wish to express our gratitude to the mathematics department at the University of Michigan for enabling this collaboration. We also thank Richard Montgomery for his helpful and constructive correspondence throughout the preparation of this paper and the referees for their useful reviews which greatly helped our exposition. This research was supported in part by NSF grant 2103026 and AFOSR grants FA9550-23-1-0215 and FA9550-23-1-0400